OpenStudy (anonymous):

Please simplify & show work! Thanks! 8 sqrt. 8-4 sqrt. 18

OpenStudy (anonymous):

\[8\sqrt{8}-4\sqrt{18}=8\sqrt{4*2}-4\sqrt{9*2}\] \[(8)(2)\sqrt{2}-(4)(3)\sqrt{2}=16\sqrt{2}-12\sqrt{2}=4\sqrt{2}\]